Net Income and OCF: During 2010, Raines Umbrella Corp. had sales of &740,000. Costs of goods...

  1. Net Income and OCF: During 2010, Raines Umbrella Corp. had sales of &740,000. Costs of goods sold, administrative and selling expenses, and depreciation expenses were $610,000, $105,000, and $140,000, respectively. In addition, the company had an interest expense of $70,000 and a tax rate of 35 percent. (Ignore any tax loss carryback or carryforward provisions.)
    1. What is Raines’ net income for 2010?
    2. What is its operating cash flow?
    3. Explain your results in (a) and (b).
Answer #1

Please provide rating..

answer a)
Computation of net income
a Sales 740,000
b Costs of goods sold, 610,000
c Administrative and selling expenses, 105,000
d Depreciation 140,000
e=a-b-c-d EBIT -115,000
f Interest 70,000
g=e-f Earning before tax -185,000
h Tax @ 35% 0
i=g-h Net income -185,000
Net income = -185,000
Answer b)
operating cash flow
EBIT = -115,000
Add : Depreciation 140,000
Operating cash flow 25,000
Answer c)
Net income include interest expenses as well as non cash expenses like depreciation
While operating cash flow only consider cash flow and ignore non cash expenses as well as interest expenses.